OpenStudy (anonymous):

1a 9 This is the problem set one A 9. : I understand the solution up until it reaches this: (2) x/2 − 1 = 1 =⇒ x = 4 and g(4) = 3f(1) − 3 = 3 (3) x/2 − 1 = 3 =⇒ x = 8 and g(8) = 3f(3) − 3 = −6 I cannot figure out why g(8) is not +6 and g(4) is not zero. Thanks for any hints!

OpenStudy (anonymous):

it is because f(3)=-1...as given in the listing of points: (-1,-1),(1,2),(3,-1), and (5,2)....therefore f((8/2)-1)=f(3)=-1 that leads to g(8)=3(-1)-3=-6

OpenStudy (anonymous):